OpenStudy (anonymous):

Determine the solution of the system. y = x + 6 y = 2x

hero (hero):

y = y x + 6 = 2x 6 = 2x - x 6 = x

OpenStudy (anonymous):

My last question!!!!

hero (hero):

You should be able to find y

OpenStudy (sirm3d):

x=6, y=12

OpenStudy (sirm3d):

x=6, y=12

OpenStudy (mayankdevnani):

let x=1 so y=1+6=7 (1,7)

OpenStudy (sirm3d):

@mayankdevnani it is a system of linear equations

OpenStudy (mayankdevnani):

yaa i know

hero (hero):

@mayandevnani, you can't just make up your own points

OpenStudy (mayankdevnani):

there are infinitely solutions @jennyjewell25

hero (hero):

There's only ONE solution mayandevnani

OpenStudy (mayankdevnani):

oh!!!!!!!

hero (hero):

For example, if you use (1,7) for y = 2x you get 7 = 2(1) 7 = 2 FALSE

OpenStudy (mayankdevnani):

no for first one @Hero

hero (hero):

The point you find has to work for BOTH equations bro. That's the whole point of systems of equations....to find the point or points that work for all given equations.

OpenStudy (mayankdevnani):

thnx...bro @Hero sorry @jennyjewell25